Acetic acid has a heat of fusion of 10.8 kJ/mol and a heat of vaporization of 24.3 kJ/mol. What is the expected value for the heat of sublimation of acetic acid?
Answer #1

AcOH CH COOH 3 AcOH(l) → AcOH(g) Hvap=24.3kJ/mol IU3 vap From hess law of heat of summation we have, U3 vap -10.8kJ/mol +24.3
Answer #2

Heat of sublimation = heat of fusion + heat of vaporization = 10.8 + 24.3 = 35.1 kJ/mol
Answer #3

Heat of sublimation = heat of fusion + heat of sublimation = 10.8 + 24.3 = 35.1 kJ/mol